Suppose you want to make an acetic acid/acetate buffer to a pH of 5.00 using 10.0 mL of 1.00 M acetic acid solution. How many milliliters of 1.00 M sodium acetate solution would you need to add? The pKa for acetate buffer is 4.75.
Answer #1

no of moles of CH3COOH = molarity * volume in L

                                           = 1*0.01   = 0.01moles

no of moles of CH3COONa   = molarity * volume in L

                                               = 1*v = v moles

PH = 5

Pka = 4.75

PH    = Pka + log[CH3COONa]/[CH3COOH]

5      = 4.75 + logv/0.01

logv/0.01   = 5-4.75

logv/0.01   = 0.25

v/0.01    = 1.7783

v       = 1.7783*0.01    = 0.017783

The volume of CH3COONa = v = 0.017783L  

                                              = 17.783ml >>>>>answer